# EarthGPT-X: Enabling MLLMs to Flexibly and Comprehensively Understand Multi-Source Remote Sensing Imagery

### v1

@@ Abstract
Recent advances in the visual-language area have developed natural multi-modal large language models (MLLMs) for spatial reasoning through visual prompting. However, due to remote sensing (RS) imagery containing abundant geospatial information that differs from natural images, it is challenging to effectively adapt natural spatial models to the RS domain. Moreover, current RS MLLMs are limited in overly narrow interpretation levels and interaction manner, hindering their applicability in real-world scenarios. To address those challenges, a spatial MLLM named EarthGPT-X is proposed, enabling a comprehensive understanding of multi-source RS imagery, such as optical, synthetic aperture radar (SAR), and infrared. EarthGPT-X offers zoom-in and zoom-out insight, and possesses flexible multi-grained interactive abilities. Moreover, EarthGPT-X unifies two types of critical spatial tasks (i.e., referring and grounding) into a visual prompting framework. To achieve these versatile capabilities, several key strategies are developed. The first is the multi-modal content integration method, which enhances the interplay between images, visual prompts, and text instructions. Subsequently, a cross-domain one-stage fusion training strategy is proposed, utilizing the large language model (LLM) as a unified interface for multi-source multi-task learning. Furthermore, by incorporating a pixel perception module, the referring and grounding tasks are seamlessly unified within a single framework. In addition, the experiments conducted demonstrate the superiority of the proposed EarthGPT-X in multi-grained tasks and its impressive flexibility in multi-modal interaction, revealing significant advancements of MLLM in the RS field.
Index Terms:
Multi-source, Spatial reasoning, Remote sensing (RS), Multi-modal large language models (MLLMs).
@@ I Introduction
Typical spatial understanding tasks in multi-modal learning include referring and grounding, which essentially require aligning regions in the image with text semantics[1]. Referring aims to accurately generate the interpretation sequence of the specific areas. Grounding requires localizing the objects or rendering object segmentation masks based on the user’s language expression[2, 3]. Most recently, enabling multi-modal large language models (MLLMs) with comprehensive spatial understanding has become the focal of research in the natural scenarios[4, 5, 6]. However, these models struggle with the remote sensing (RS) imagery, which features multi-source, multi-viewpoint, and scale variations due to different imaging conditions[7]. Although current RS MLLMs can handle a wide range of RS visual tasks[8, 9], they primarily focus on holistic image understanding, lacking spatial perception. They solely rely on text instructions which can lead to ambiguity and limited context. It is indispensable for establishing a comprehensive and spatial-aware MLLM to interpret RS images under flexible interactions.
@@
Notably, in the human world, people can achieve an all-around understanding of visual information, allowing for both zoom-in fine-grained analyses and zoom-out coarse-grained overviews[10]. Inspired by this, universal MLLM and multi-grained prompting learning have been extensively studied in the visual-language learning community[11]. Remarkable advances of MLLMs include close-source Open-AI’s GPT series[12, 13, 14], open-source Meta’s LLaMA series[15, 16], and parameter-efficient LLaMA-Adapter series[17, 18], which demonstrate versatile capabilities of handling multi-modal various visual tasks. On top of that, visual prompting models advance the field by injecting visual prompts beyond text instructions to enable multi-grained visual reasoning and enhance spatial perception[19]. Specifically, representative works comprise promptable segmentation model segment anything (SAM)[20], region-of-interest perception model GPT4RoI [4], refer-and-ground MLLM Ferret[1], etc. However, the aforementioned models are tailored for the natural imagery, it is challenging to directly adapt them to the diverse RS data.
@@
Figure 1: An overview of the proposed flexible and comprehensive spatial understanding framework for RS imagery: (a) Task: various multi-grained referring and grounding tasks; (b) Model: the proposed model consists of an image encoder, a visual prompts encoder, a text tokenizer, a LLM, and a pixel perception module; (c) Data: the training dataset M-RSVP is transformed based on existing multi-source RS data including optical, SAR, and infrared.
@@
Limited efforts in the RS domain have been made to enhance spatial understanding. A key example is RSPrompter [21], which is based on SAM and develops an automated prompts generation approach to facilitate interactive segmentation. However, RSPrompter is only designed for instance segmentation. RSVG [2] defines the visual grounding tasks in the RS domain and can locate the referred objects using boxes under natural language expression. Recent notable works, such as EarthGPT[8] and Geochat[9], have realized RS holistic and regional imagery comprehension under the unified visual-language learning framework, facilitating the development of MLLMs in the RS domain. Nevertheless, RSVG, EarthGPT, and Geochat solely support language interaction without visual prompts to realize referring expression comprehension. Notably, EarthMarker[22] is constructed as the first visual prompting learning framework in the RS domain and is capable of region-level and point-level RS imagery interpretation. Regretfully, EarthMarker is mainly oriented towards optical images and lacks support for multi-source imagery and free-form visual prompts, hampering flexible human-AI interaction and comprehensive understanding of RS data. Thus, the study on enabling MLLMs to flexibly interpret multi-source RS images at any granularity remains in its infancy stages.
@@
To resolve the limitations of existing models and draw inspiration from how humans understand multi-modal, multi-grained information through flexible interactions, a spatial MLLM named EarthGPT-X is proposed based on visual prompt learning. EarthGPT-X is designed for a comprehensive understanding of multi-source RS imagery under flexible visual prompts (e.g., points, boxes, circles, and scribbles), and unifying referring and grounding tasks within a free-form region-text alignment architecture. As depicted in Fig. 1, the EarthGPT-X architecture consists of three interconnected components: multi-grained RS image interpretation tasks, a visual prompting model, a construction pipeline of training datasets. Specifically, EarthGPT-X not only allows for zoom-in fine-grained reasoning (e.g., free-form region captioning, referring object classification, and object inter-relationship analyses), but also supports zoom-out coarse-grained tasks (e.g., image-level captioning, scene classification). Secondly, the key of the model is to realize multi-modal content integration including images, visual prompts, and text instructions. The multi-modal signals are then fed into the large language models (LLM) for subsequent mixed tuning. Additionally, to unify referring and grounding in a single framework, the pixel perception module is designed to generate binary segmentation masks related to referring objects. Thirdly, the training dataset named M-RSVP is constructed, transforming from existing multi-source datasets, which is uniformly reorganized and relabeled as image-visual prompts-text triples, with a portion re-annotated using GPT-4V[23] to support the generation of longer and more detailed descriptions[22].
In addition, observed that existing synthetic aperture radar (SAR) and infrared datasets are limited to object detection tasks, which are insufficient for developing complex visual reasoning capabilities of MLLMs. To address this shortcoming, the multi-domain data are integrated for generalizing general-domain knowledge into SAR and infrared scenarios. Furthermore, to allow for free-form visual prompts data, a noise augmentation method is proposed to add noise that either partially covers the ground truth or extends beyond it to simulate free-form shapes. Based on the newly constructed training datasets, a one-stage all-in-one training strategy is proposed for mixed tuning, facilitating cross-domain data mutual learning and alleviating the need for manual effort for manually setting during the training.
@@
To further investigate EarthGPT-X’s superior comprehensive interpretation ability on multi-source imagery, we compare it with recent state-of-the-art (SOTA) natural and RS MLLMs. Impressively, EarthGPT-X can realize multi-source data reasoning accurately and provide a more comprehensive and detailed understanding of RS images. Furthermore, to evaluate the effectiveness of the EarthGPT-X’s structure, we conduct ablation studies to assess the designed components’ contributions to the overall performance. Experimental results demonstrate that the proposed EarthGPT-X not only achieves competitive performances on vision-language tasks, but also creates impressive results on referring and grounding tasks.
@@
  • A RS Multi-source Spatial MLLM, EarthGPT-X. A pioneering MLLM named EarthGPT-X is proposed, specifically designed for a comprehensive spatial understanding of multi-source RS imagery. EarthGPT-X advances existing RS MLLMs by enabling flexible multi-grained imagery interpretation, and offers zoom-in and zoom-out insight, powered by designed free-form visual prompts. Furthermore, EarthGPT-X unifies referring and grounding tasks into a visual prompting learning framework, facilitates an adaptable understanding of RS data, and represents a significant milestone in developing versatile MLLMs on RS scenarios.
  • First Multi-source Visual Instruction Dataset, M-RSVP. A multi-source multi-modal visual-prompt-based instruction dataset named M-RSVP, comprising over 200 k image-visual prompts-text triples, is constructed. M-RSVP fills the gap of multi-source visual prompting data in the RS domain, facilitating a multi-grained understanding of RS imagery across various scenes, diverse sources, and varying imaging conditions.
  • Flexible Multi-modal Content Interaction. A multi-modal content integration mechanism is developed, enhancing the mutual understanding between dense image representations, sparse visual prompt features, and text instruction embeddings. This mechanism leads to a deeper comprehension of the referring regions in holistic images. Moreover, by combining with a noise augmentation method for simulating free-form shapes, this mechanism endows the proposed EarthGPT-X with comprehensive RS imagery understanding under flexible human-AI interaction.
  • Cross-domain Knowledge Integration. A cross-domain one-stage fusion training strategy is proposed, transferring and generalizing universal domain-level knowledge into SAR and infrared scenarios. This approach effectively mitigates the limitations of existing SAR and infrared data interpretation, which are typically restricted to object detection tasks and hinder the potential to perform complex multi-task visual reasoning. Furthermore, the one-stage all-in-one training facilitates cross-domain data mutual perception and reduces the reliance on extensive manual effort.
@@ II Related Work
II-A LLMs and MLLMs in Natural Scenarios
The rapid development of LLMs has significantly advanced natural language processing (NLP), demonstrating remarkable capabilities in diverse contexts. Specifically, GPT-3 [12] has demonstrated the power of large-scale models, containing billions to trillions of parameters. InstructGPT [24] and ChatGPT [13] have further enhanced this by showcasing remarkable fluency and adaptability in various conversational tasks. Impressively, the open-source community has significantly contributed to advancing LLMs, with resources like LLaMA series [25] enhancing the instruction-following capabilities of these models. Recent efforts, including Alpaca [26], Vicuna [27], and GPT-4-LLM [28], have introduced full fine-tuning techniques to effectively equip LLMs with robust instruction-following abilities. Additionally, LLaMA-Adapter [18] has demonstrated that parameter-efficient fine-tuning can achieve comparable results with full parameter fine-tuning, offering a more efficient approach to model adaptation.
@@ II-A LLMs and MLLMs in Natural Scenarios
Notably, an essential step toward achieving general intelligence is integrating multi-modal perception into LLMs, allowing them to handle multi-modal data beyond text. Several works [29, 30, 31, 32] focus on incorporating multi-modal inputs to support task-specific objectives across various modalities. Models like VisualGPT [29], BLIP [30], Flamingo [32], and Kosmos-1 [33] have shown strong multi-modal reasoning abilities by aligning LLMs with visual data. This alignment is achieved using techniques like projection layers, zero-shot attention mechanisms, and intermediate networks, as demonstrated by MiniGPT-4 [34], LLAMA-Adapter V2 [17], and mPLUG-Owl [35]. In addition, with multi-modal LLMs continuing to evolve, other modalities include audio, video [36, 37], and 3D point clouds [38, 39, 40] are being integrated into MLLMs.
II-B MLLMs in RS Scenarios
Inspired by MLLMs in natural domains, MLLMs in the RS field have also made significant strides. Several pioneering RS MLLMs have emerged recently. For example, RSGPT [41], which fine-tunes Instruct-BLIP [42] on a high-quality image-text pairs dataset, demonstrates strong capabilities in image captioning and visual question answer (VQA). However, RSGPT struggles with other visual tasks including object classification and detection, lacking generalization abilities. A key advance is Geochat [9], which is the first versatile MLLM designed to solve multiple tasks for optical RS images. EarthGPT [8] introduces a universal MLLM capable of handling multi-source RS imagery and a broad range of RS visual tasks, and supports regional visual reasoning. In addition, another RS MLLM called LHRS-Bot[43] has established the LHRS-Bench to assist the RS community in evaluating RS-domain MLLMs across diverse evaluation metrics. These models have undoubtedly advanced the development of MLLMs within the RS-specific domain. Nevertheless, EarthGPT, Geochat, and LHRS-Bot solely allow language interaction in multi-turn conversations to handle visual tasks, and fail to support referring comprehension under visual prompt instruction, which restricts flexible interaction in real-world applications.
@@ II-C Prompting Learning
Prompt learning has emerged as a key area of research in NLP [12]. Notable works such as AutoPrompt [44] and CoOp [45] focus on automating prompt template generation for language and vision-language models, reducing the need for manual crafting. In parallel, language prompting has been applied to open-vocabulary detection models, including DetPro [46] and PromptDet [47]. While language prompting is well-developed, visual prompting remains underexplored. A significant advancement in this area is the SAM [20] model, which supports various segmentation prompts to improve zero-shot performance. However, SAM lacks object semantic labels, which is addressed by Semantic-SAM [48]. Moreover, models like GPT4RoI [4] and RegionBlip [5] leverage spatial boxes and combine language with region-of-interest input for regional recognition. Colorful Prompting Tuning (CPT) [49] uses color-based prompts to enhance MLLMs performance. Osprey [3] goes further by incorporating fine-grained mask regions into language instructions for pixel-level understanding. Ferret [1] is capable of understanding spatial referring of any shape or granularity within an image and accurately grounding open-vocabulary descriptions. Additionally, SPHINX-V [50] supports free-form visual prompts to achieve pixel-level image comprehension. Mentioned above models are designed for the natural domain.
Note that EarthMarker [22] stands as the pioneering visual prompting framework tailored for the RS domain, specifically designed to interpret RS imagery at both region and point levels. EarthMarker constructs a large-scale RS visual prompting dataset and leverages visual prompts learning to endow MLLM with region-level and point-level RS imagery interpretation capabilities. However, EarthMarker is primarily focused on optical images and lacks support for multi-source imagery and free-form visual prompts. This limitation significantly hinders its potential for flexible interactions and restricts its ability to provide a comprehensive understanding of the diverse and complex characteristics of RS data and broader applicability in real-world RS scenarios.
In summary, existing RS MLLMs have constructed the visual-language learning diagram to comprehend RS imagery and support language interaction in multi-turn dialogues. More advanced than existing MLLMs and EarthMarker, the proposed EarthGPT-X is designed for multi-source imagery comprehension using a dual-prompt system of text instructions and visual prompts. With free-form visual prompting, EarthGPT-X offers comprehensive and flexible visual perception. A straightforward side-by-side comparison of EarthGPT-X versus the recent RS MLLMs is shown in Tab. I.
TABLE I: Comparisons of EarthGPT-X v.s. Recent MLLMs in RS domain.
Model Text Image Visual Prompts Pixel-level
Instruction Optical Multi-source Point Box Free-form Grounding
RSGPT[41]
RS-LLaVA[51]
GeoChat[9]
EarthGPT[8]
Popeye[52]
SkyEyeGPT[53]
EarthMarker[22]
EarthGPT-X

### v2

# EarthGPT-X: A Spatial MLLM for Multi-level Multi-Source Remote Sensing Imagery Understanding with Visual Prompting

@@ Abstract
Recent advances in natural-domain multi-modal large language models (MLLMs) have demonstrated effective spatial reasoning through visual and textual prompting. However, their direct transfer to remote sensing (RS) is hindered by heterogeneous sensing physics, diverse modalities, and unique spatial scales. Existing RS MLLMs are mainly limited to optical imagery and plain language interaction, preventing flexible and scalable real-world applications. In this article, EarthGPT-X is proposed, the first flexible spatial MLLM that unifies multi-source RS imagery comprehension and accomplishes both coarse-grained and fine-grained visual tasks under diverse visual prompts in a single framework. Distinct from prior models, EarthGPT-X introduces: 1) a dual-prompt mechanism combining text instructions with various visual prompts (i.e., point, box, and free-form) to mimic the versatility of referring in human life; 2) a comprehensive multi-source multi-level prompting dataset, the model advances beyond holistic image understanding to support hierarchical spatial reasoning, including scene-level understanding and fine-grained object attributes and relational analysis; 3) a cross-domain one-stage fusion training strategy, enabling efficient and consistent alignment across modalities and tasks. Extensive experiments demonstrate that EarthGPT-X substantially outperforms prior nature and RS MLLMs, establishing the first framework capable of multi-source, multi-task, and multi-level interpretation using visual prompting in RS scenarios. The code and dataset are available at https://github.com/wivizhang/EarthGPT-X.
@@ I Introduction
Typical spatial understanding tasks in multi-modal learning require aligning image regions with text semantics[1]. Beyond holistic description, such tasks involve accurately interpreting specific regions or objects based on user instructions[2, 3]. Most recently, enabling multi-modal large language models (MLLMs) with comprehensive spatial understanding has become the focus of research in natural scenarios[4, 5, 6]. However, these models struggle with the remote sensing (RS) imagery, which features multi-source, multi-viewpoint, and scale variations due to different imaging conditions[7]. Although existing RS MLLMs can handle multiple RS visual tasks[8, 9], they primarily focus on holistic image understanding, lacking spatial perception. They solely rely on text instructions, which can lead to ambiguity and limited context. It is indispensable to develop a spatial-aware MLLM to interpret RS imagery under flexible interactions.
@@
Notably, in the human world, people can achieve an all-around understanding of visual information, allowing for both zoom-in fine-grained analyses and zoom-out coarse-grained overviews[10]. Inspired by this, visual prompting models advance the field by injecting visual prompts beyond text instructions to enable multi-grained visual reasoning and enhance spatial perception[11]. Representative works include the promptable segmentation model Segment Anything (SAM) [12], which enables class-agnostic delineation from visual or text prompts and shows strong generalization as a versatile backbone. The region-of-interest model GPT4RoI [4] integrates large language models (LLMs) for localized reasoning, while Ferret [1] focuses on fine-grained grounding, which can accept diverse region inputs, such as points and bounding boxes. Additionally, Wu et al. [13] provide a review of visual prompting approaches, including visual prompting, prompt generation, and prompt learning. However, designed for natural images, these models cannot be directly adapted to the heterogeneous RS domain due to fundamental differences in sensing physics, modalities, and scales.
@@
Figure 1: Overview of the proposed framework for flexible and comprehensive spatial understanding in RS imagery: (a) Tasks: a wide range of multi-granularity visual tasks; (b) Model: EarthGPT-X integrates an image encoder, a visual prompt encoder, a text tokenizer, an LLM, and a pixel perception module; (c) Data: the training dataset M-RSVP is constructed from multi-source RS data (optical, SAR, infrared) and further enriched with GPT-4V–based annotations.
@@
Limited efforts in the RS domain have been made to enhance spatial understanding. A key example is RSPrompter [14], which is based on SAM and develops an automated prompts generation approach to facilitate interactive segmentation. However, RSPrompter is only designed for instance segmentation. RSVG [2] defines the visual grounding tasks in the RS domain and can locate the referred objects using boxes under natural language expressions. Recent notable works, such as EarthGPT[8] and Geochat[9], have realized RS holistic and regional imagery comprehension under the unified visual-language learning framework, facilitating the development of MLLMs in the RS domain. Nevertheless, these models solely support language interaction without visual prompts to realize referring expression comprehension. Notably, EarthMarker[15] is a visual prompting model in the RS domain, however, it lacks support for multi-source imagery and free-form visual prompts, hampering flexible human-AI interaction and comprehensive understanding of RS data. Thus, enabling MLLMs to interpret multi-source RS imagery at any granularity remains in its infancy.
@@
To overcome these limitations, we propose EarthGPT-X, a spatial MLLM built upon visual prompt learning, tailored for multi-source RS imagery, supporting diverse visual prompt types (e.g., points, boxes, circles, and scribbles), and enabling comprehensive interpretation from overall scene understanding to fine-grained regional analysis, thereby offering more flexible and scalable visual-language interaction than prior approaches. As depicted in Fig. 1, the EarthGPT-X architecture consists of three interconnected components: multi-grained RS image interpretation tasks, a visual prompting model, and a construction pipeline of training datasets. Specifically, EarthGPT-X not only allows for zoom-in fine-grained reasoning (e.g., free-form region captioning, referring object classification, and object inter-relationship analyses), but also supports zoom-out coarse-grained tasks (e.g., image-level captioning, scene classification). Secondly, the key of the model is to realize multi-modal content integration, including images, visual prompts, and text instructions. The multi-modal signals are then fed into the LLMs for a one-stage all-in-one training. Additionally, a pixel perception module is integrated to extend the model with object segmentation masks. Most importantly, the training dataset named M-RSVP is constructed, transforming from existing multi-sensor datasets, which are uniformly reorganized and relabeled as image-visual prompts-text triples. In addition, a portion of the data re-annotated using GPT-4V[16] to generate longer and more detailed descriptions. A side-by-side comparison of EarthGPT-X with recent RS MLLMs is presented in Tab. I, highlighting its advancement in modality coverage, dataset quality, prompting flexibility, and fine-grained multi-task reasoning.
@@
To further validate EarthGPT-X’s comprehensive interpretation ability on multi-source imagery, we compare it with recent state-of-the-art (SOTA) natural and RS MLLMs. The results indicate that EarthGPT-X achieves accurate multi-source reasoning and provides more detailed RS image understanding. Overall, EarthGPT-X not only attains competitive performance on vision–language benchmarks, but also demonstrates remarkable advantages in complex object-level reasoning tasks.
@@
  • EarthGPT-X: First Multi-source Spatial MLLM. We propose EarthGPT-X, a spatial MLLM designed for multi-source (i.e., optical, SAR, and infrared) RS imagery understanding. It advances existing RS MLLMs by supporting free-form visual prompts (e.g., points, boxes, scribbles) to enable flexible multi-granularity interaction. EarthGPT-X unifies diverse spatial reasoning tasks into a visual prompting framework, providing a generalizable interpretation across heterogeneous RS data.
  • M-RSVP: First Multi-source Visual Prompting Dataset. The first multi-source visual prompting dataset for RS, titled M-RSVP, consisting of over 650k image-visual prompt-text triples, is constructed. The constructed M-RSVP bridges the gap of multi-source visual prompting data in the RS domain, enabling multi-granularity interpretation across diverse scenes and sensors.
  • Multi-domain Multi-modal Multi-level Interaction. A hybrid signal mutual understanding mechanism is designed to strengthen the interplay between dense image features, sparse visual prompts, and text instructions, thereby improving the interpretation of referring regions within holistic images. Moreover, our cross-domain one-stage fusion training significantly streamlines the training procedure of MLLMs, while ensuring seamless integration of multi-source knowledge.
  • Superior Performance. Extensive experiments benchmark EarthGPT-X against SOTA natural and RS MLLMs across diverse multi-level visual tasks and multi-source imagery. Furthermore, EarthGPT-X also excels in complex reasoning tasks, such as object inter-relationship analysis. These advances collectively represent a significant milestone toward developing versatile and flexible MLLMs for RS scenarios.
TABLE I: Comparisons of EarthGPT-X v.s. Recent MLLMs.
Model Textual Modalities RS Dataset Tasks Visual Prompts Pix.
Prompt Opt. Multi. Pub. Refi. Regu. Comp. Point Box Free-form Grd.
Shikra (Arxiv 2023)[6]
GPT4RoI (ECCV 2024)[17]
ViP-LLaVA (CVPR 2024)[18]
RS-LLaVA (RS 2024)[19]
GeoChat (CVPR 2024)[9]
EarthGPT (TGRS 2024)[8]
Popeye (JSTAR 2024)[20]
LHRS-Bot (ECCV 2024)[21]
VHM (AAAI 2025)[22]
EarthDial (CVPR 2025)[23]
EarthMarker (TGRS 2025)[15]
EarthGPT-X (Ours)
Note: Comparisons in terms of modalities (optical or multi-source), dataset construction (converted directly from public data or Refined using powerful MLLMs such as GPT-4V or Gemini), supported tasks (Regular visual tasks or complex reasoning such as inter-relationship analysis), visual prompting capabilities, and extended pixel-level grounding.
@@ II Related Work
II-A MLLMs in Natural Scenarios
@@
The rapid development of LLMs has significantly advanced natural language processing (NLP), demonstrating remarkable capabilities in diverse contexts [24, 25, 26]. Then, an essential step toward achieving general intelligence is integrating multi-modal perception into LLMs, allowing them to handle multi-modal data beyond text. Several works [27, 28, 29, 30] focus on incorporating multi-modal inputs to support task-specific objectives across various modalities. Models like VisualGPT [27], Flamingo [30], and SPHINX-X[31] have shown strong multi-modal reasoning abilities by aligning LLMs with visual data. This alignment is achieved through projection layers, zero-shot attention, and intermediate networks, as in MiniGPT-4 [32], LLAMA-Adapter V2 [33], and mPLUG-Owl [34]. More recent advances include Qwen2.5-VL [35], MiniCPM-V2.6 [36], InternVL2.5 [37], InternLM-XComposer2.5 [38], and DeepSeek-VL2 [39]. These enhanced versions advance prior frameworks by broadening input modalities, leveraging higher-quality large-scale datasets, and extending the scope of multi-modal reasoning tasks.
II-B Specialist and MLLMs in RS Scenarios
RS imagery encompasses diverse modalities such as SAR, optical, and infrared, each capturing complementary physical characteristics of the observed scene. Recent advanced SAR detection adopts generative or predictive paradigms, e.g., MaDiNet [40] with Gamma Diffusion–based bounding-box generation and SAR-JEPA [41] with joint-embedding predictive recognition. In optical understanding, specialist models such as CSDS [42] are designed for scene classification, while Qonly [43] targets VQA tasks. For infrared detection, IRSAM [44] enhances SAM to achieve more discriminative representations. Although effective for their respective tasks, these specialist models lack the ability to perform multiple visual tasks in a unified architecture, thus limiting scalability and hindering broader real-world deployment.
Inspired by the success of MLLMs in natural image domains, RS research has also begun to adopt MLLMs with notable progress. RSGPT [45], built on Instruct-BLIP [46], demonstrates strong performance in captioning and VQA but struggles with classification and detection. Cross-modal transfer learning in the RS MLLM field has recently gained traction, with representative works exemplifying different stages of progress [47, 48]. GeoChat [9] represents the first versatile RS MLLM for multiple optical tasks, while SkyEyeGPT [21] further extends broader task coverage, EarthGPT [8] bridging MLLMs with optical, SAR, and infrared domains. LHRS-Bot [49] additionally contributes by introducing LHRS-Bench for standardized evaluation. These efforts are collectively paving the way for more unified and versatile frameworks. However, they remain limited to text-only prompts, lacking flexible interactive comprehension under visual instructions.
@@ II-C Prompting Learning
Prompt learning has become a key research direction in NLP [24], with works such as AutoPrompt [50] and CoOp [51] automating prompt design for language and vision–language models. Subsequently, the idea evolved into visual prompting [13], marking a broader shift from language-based to multi-modal prompts. A milestone in this line is the Segment Anything Model (SAM) [12], which enables class-agnostic segmentation via point, box, or text prompts and demonstrates strong generalization across diverse domains. GPT4RoI [4] and RegionBlip [5] leverage spatial boxes and combine language with region-of-interest input for regional recognition. Colorful Prompting Tuning (CPT) [52] uses color-based prompts to enhance MLLMs’ performance. Osprey [3] goes further by incorporating fine-grained mask regions into language instructions for pixel-level understanding. ViP-LLava[47] supports arbitrary visual cues like scribbles and arrows to interpret images flexibly. Wu et al. [13] provide a comprehensive survey of visual prompting approaches, covering visual prompting, prompt generation, and prompt learning. These models are primarily designed for the natural image domain and fail to generalize to the RS multi-source data. Notably, EarthMarker [15] is a visual prompting framework tailored for the RS data, but it remains confined to optical imagery and does not support flexible prompt types. These gaps highlight the necessity of a specialized model that can support multi-source imagery and flexible interaction in RS scenarios, motivating the development of the proposed model.